"Question1\n\nGiven A= [\u25a0(2&2&1@1&3&1@1&2&2)] \n\n\n\n\n1.1 find the eigenvalues of A\n\n1.2. Determine an eigenvector for the eigenvalue = 5\n\n\n\nQuestion2\nA periodic function f(x) with period 2 \u03c0 is defined by :\n f(x)= [ (x+\u03c0)\/2 -\u03c0\u02c2 X\u02c2 0, (x-\u03c0)\/2 0\u02c2 X\u02c2 \u03c0]\n\nDetermine the Fourier series for the periodic function"


1
Expert's answer
2021-08-06T11:35:48-0400

1.


"A=\begin{pmatrix}
 2 & 2 & 1 \\
 1 & 3 & 1 \\
 1 & 2 & 2
\end{pmatrix}"

"\det(A-\lambda I)=\begin{vmatrix}
 2-\lambda & 2 & 1 \\
 1 & 3-\lambda & 1 \\
 1 & 2 & 2-\lambda \\
\end{vmatrix}=0"

"(2-\lambda)^2(3-\lambda)+2+2-3+\lambda-4+2\lambda-4+2\lambda=0"

"(4-4\lambda+\lambda^2)(3-\lambda)+5\lambda-7=0"

"12-4\lambda-12\lambda+4\lambda^2+3\lambda^2-\lambda^3+5\lambda-7=0"

"-\lambda^3+7\lambda^2-11\lambda+5=0"

"-\lambda^2(\lambda-1)+6\lambda(\lambda-1)-5(\lambda-1)=0"


"-(\lambda-1)(\lambda^2-6\lambda+5)=0"

"-(\lambda-1)^2(\lambda-5)=0"

The roots are "\lambda_1=1, \lambda_2=1, \lambda_3=5."

These are the eigenvalues.

"\lambda=5"


"\begin{pmatrix}
 2-\lambda & 2 & 1 \\
 1 & 3-\lambda & 1 \\
 1 & 2 & 2-\lambda \\
\end{pmatrix}=\begin{pmatrix}
 -3 & 2 & 1 \\
 1 & -2 & 1 \\
 1 & 2 & -3
\end{pmatrix}"

"R_1=R_1\/(-3)"


"\begin{pmatrix}
 1 & -2/3 & -1/3 \\
 1 & -2 & 1 \\
 1 & 2 & -3
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 1 & -2/3 & -1/3 \\
 0 & -4/3 & 4/3 \\
 1 & 2 & -3
\end{pmatrix}"

"R_3=R_3-R_1"


"\begin{pmatrix}
 1 & -2/3 & -1/3 \\
 0 & -4/3 & 4/3 \\
 0 & 8/3 & -8/3
\end{pmatrix}"

"R_2=-3R_2\/4"


"\begin{pmatrix}
 1 & -2/3 & -1/3 \\
 0 & 1 & -1 \\
 0 & 8/3 & -8/3
\end{pmatrix}"

"R_1=R_1+2R_2\/3"


"\begin{pmatrix}
 1 & 0 & -1\\
 0 & 1 & -1 \\
 0 & 8/3 & -8/3
\end{pmatrix}"

"R_3=R_3-8R_2\/3"


"\begin{pmatrix}
 1 & 0 & -1\\
 0 & 1 & -1 \\
 0 & 0 & 0
\end{pmatrix}"

"\begin{pmatrix}
 1 & 0 & -1\\
 0 & 1 & -1 \\
 0 & 0 & 0
\end{pmatrix}\begin{pmatrix}
 x_1\\
 x_2 \\
 x_3
\end{pmatrix}=\begin{pmatrix}
 0\\
 0 \\
 0
\end{pmatrix}"

If we take "x_3=t," then "x_1=t, x_2=t."


"\vec x=\begin{pmatrix}
 t\\
 t\\
 t
\end{pmatrix}=\begin{pmatrix}
 1\\
 1\\
 1
\end{pmatrix}t""\vec v=\begin{pmatrix}
 1\\
 1\\
 1
\end{pmatrix}"

This is the eigenvector.



"\lambda=1"

"\begin{pmatrix}
 2-\lambda & 2 & 1 \\
 1 & 3-\lambda & 1 \\
 1 & 2 & 2-\lambda \\
\end{pmatrix}=\begin{pmatrix}
 1 & 2 & 1 \\
 1 & 2 & 1 \\
 1 & 2 & 1
\end{pmatrix}"

"R_2=R_2-R_1"

"\begin{pmatrix}
 1 & 2 & 1 \\
 0 & 0 & 0 \\
 1 & 2 & 1
\end{pmatrix}"

"R_3=R_3-R_1"

"\begin{pmatrix}
 1 & 2 & 1 \\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{pmatrix}"


"\begin{pmatrix}
 1 & 2 & 1\\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{pmatrix}\begin{pmatrix}
 x_1\\
 x_2 \\
 x_3
\end{pmatrix}=\begin{pmatrix}
 0\\
 0 \\
 0
\end{pmatrix}"

If we take "x_2=t, x_3=s," then "x_1=-2t-s."

"\vec x=\begin{pmatrix}
 -2t-s\\
 t\\
 s
\end{pmatrix}=\begin{pmatrix}
 -2\\
 1\\
 0
\end{pmatrix}t+\begin{pmatrix}
 -1\\
 0\\
 1
\end{pmatrix}s"

"\vec u=\begin{pmatrix}
 -2\\
 1\\
 0
\end{pmatrix}, \vec w=\begin{pmatrix}
 -1\\
 0\\
 1
\end{pmatrix}"

These are the eigenvectors.


2.


"a_0=\dfrac{2}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)dx"

"=\dfrac{2}{\pi}\displaystyle\int_{-\pi}^{0}\dfrac{x+\pi}{2}dx+\dfrac{2}{\pi}\displaystyle\int_{0}^{\pi}\dfrac{x-\pi}{2}dx"

"=\dfrac{2}{\pi}[\dfrac{x^2}{4}+\dfrac{\pi x}{2}]\begin{matrix}
 0 \\
 -\pi & d
\end{matrix}+\dfrac{2}{\pi}[\dfrac{x^2}{4}-\dfrac{\pi x}{2}]\begin{matrix}
 \pi \\
 0
\end{matrix}"

"=\dfrac{\pi}{2}-\dfrac{\pi}{2}=0"



"a_n=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)\cos{nx}dx"

"=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{0}\dfrac{x+\pi}{2}\cos{nx}dx+\dfrac{1}{\pi}\displaystyle\int_{0}^{\pi}\dfrac{x-\pi}{2}\cos{nx}dx"

"=-\dfrac{\cos{\pi n}-1}{2\pi n^2}+\dfrac{\cos{\pi n}-1}{2\pi n^2}=0"



"b_n=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)\sin{nx}dx"

"=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{0}\dfrac{x+\pi}{2}\sin{nx}dx+\dfrac{1}{\pi}\displaystyle\int_{0}^{\pi}\dfrac{x-\pi}{2}\sin{nx}dx"

"=-\dfrac{1}{2n}-\dfrac{1}{2n}=-\dfrac{1}{n}"




"f(x)=-\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n}\sin{nx}"